# Preliminary Ranking of WMT25 General Machine Translation Systems

### v1

@@ Introduction
We present the preliminary ranking of the WMT25 General Machine Translation Shared Task, in which MT systems have been evaluated using automatic metrics. As this ranking is based on automatic evaluations, it may be biased in favor of systems that employ re-ranking techniques, such as Quality Estimation re-ranking or Minimum Bayes Risk decoding. The official WMT25 ranking will be based on human evaluation, which is more reliable and will supersede the automatic ranking.
The purpose of this report is not to present the final findings of the General MT task, but rather to share preliminary results with task participants, which may be useful when preparing their system submission papers.
@@
  • Constrained systems are those using only publicly available training data and models. The maximum size of their parameter counts is 20B and participants are required to release their weights under an open license.
  • Unconstrained systems (marked with gray) are all the remaining systems, with no limitations on their training data, model sizes and no requirements to publish their model weights. Systems where the relevant information is not publicly known also fall into the unconstrained category.
@@
Details of all systems are going to be available in the upcoming WMT25 findings. In addition to participating systems, we also collect open-weight and proprietary LLMs together with three popular commercial MT systems. For each provider, we selected their largest/best performing model for each of the subtracks (when applicable).
Constrained systems: AyaExpanse-8B, CommandR7B, EuroLLM-9B, Gemma-3-12B, Llama-3.1-8B, Mistral-7B, NLLB, Qwen2.5-7B, TowerPlus-9B
Unconstrained systems: AyaExpanse-32B, Claude-4, CommandA, DeepSeek-V3, EuroLLM-22B, Gemma-3-27B, Gemini-2.5-Pro, GPT-4.1, Llama-4-Maverick, Mistral-Medium, ONLINE-B, ONLINE-G, ONLINE-W, Qwen3-235B, TowerPlus-72B
We used a zero shot instruction following approach, translating data on a document-level whenever possible, having a paragraph-level backup for failed translations. The instructions were provided in the blindset. This may hurt some of the systems trained for specific MT instructions such as TowerLLM or EuroLLM, we mark them with [M].
We turned off the reasoning for Qwen3-235B, however, we didn’t set the reasoning budget for Gemini-2.5-Pro which increased output tokens count 6.6 times making it the most expensive model in the evaluation.
The code for collecting translations is available at github.com/wmt-conference/wmt-collect-translations and we marked all systems collected by us with $\blacktriangle$ .
Evaluated Data
We evaluated 32 language pairs: half of them will be evaluated by humans, while the other half belong to the multilingual subtrack and will rely solely on automatic ranking.
Most language pairs are in the English-to-X direction and contain approximately 37k words. Each segment contains about 100 words, which typically corresponds to paragraphs, although sometimes natural paragraphs had to be further segmented to meet this constraint. The data are aggregated into documents. The test sets combine material from four domains:
@@
Participants could use image and video modalities (when available) to improve their translations; however, their use was not required. Humans providing reference translations were offered these extra modalities, too. Language pairs with a non-English source have a similar distribution but differ slightly in domains and sizes.
We do not provide sentence splitting; consequently, many segments contain multiple sentences.
We release all data, including references, system outputs, automatic segment scores, or latex sources of this document at: github.com/wmt-conference/wmt25-general-mt.
@@
We call the automatic ranking described in the rest of this section “AutoRank”. Compared to last year, both the set of automatic metrics and the aggregation procedure changed slightly.
Metrics used.
For each language pair (except where noted below), we combine three families of evaluation methods:
  • LLM-as-a-Judge (reference-less). GEMBA-ESA (Kocmi and Federmann, 2023) with two independent judges: GPT-4.1 (OpenAI, 2025) and Command A (Cohere Team, 2025), both used in a reference-less setting.
  • Trained reference-based metrics. Two reference-based supervised metrics explicitly trained to approximate human judgments of translation quality: MetricX-24-Hybrid-XL (Juraska et al., 2024) and XCOMET-XL (Guerreiro et al., 2024).
  • Trained Quality Estimation (QE). One QE metric trained to mimic human judgments without a reference: CometKiwi-XL (Rei et al., 2023).
Including both reference-based and reference-less (or QE) methods balances complementary failure modes: reference-based metrics typically achieve higher correlation with human judgments when references are high-quality, whereas reference-less methods reduce susceptibility to reference bias when references are suboptimal (Freitag et al., 2023). A known pitfall for multilingual QE is that it can be fooled by fluent output in the wrong target language; in contrast, the GEMBA-ESA prompt explicitly specify the target language, which should mitigate this issue.
The use of LLM-as-a-judge metrics (GEMBA-ESA) is intended to mitigate biases by models employing re-ranking or similar techniques during training or inference. Nevertheless, some systems incorporated GEMBA directly as their reward model.
For each metric and language pair, the system-level score of an MT system is computed as the average of the metric’s paragraph-level (segment-level) scores over all translations the system produced on the test set for that language pair. For language pairs without human references, we exclude CometKiwi-XL from the corresponding AutoRank computation, since MetricX-24-Hybrid-XL and XCOMET-XL are hybrid metrics and can be run in reference-less (QE) mode, thus already providing the QE signal from trained metrics for those pairs.
@@
For the two most low-resource target languages, i.e., Bhojpuri and Maasai, we rely solely on chrF++ (Popović, 2017) because it is not known whether the above metrics are reliable in these settings (Falcão et al., 2024; Singh et al., 2024; Wang et al., 2024; Sindhujan et al., 2025) and human references are available. We compute chrF++ using sacrebleu (Post, 2018).
From system-level scores to AutoRank
@@
This year, we received 36 unique teams, [^1] We received 43 different teams, however, 7 of them have withdrew or been disqualified the highest number of participants ever. We are not able to evaluate them all with human annotators. Therefore, we select a subset of about 18 systems per language pair (some language pairs have this system count higher) which will be evaluated by humans with the Error Span Annotation protocol (Kocmi et al., 2024). For the remaining systems, AutoRank is going to be the official final ranking.
When selecting systems for human evaluation, we prioritize constrained systems over unconstrained systems. Therefore, we select the systems for human evaluation based on the following two rules:
@@
    We select top eight constrained systems ignoring unconstrained systems.
@@
    Then, we take the top performing systems until we have total of 18 systems selected for human evaluation.
@@
A key limitation of our evaluation is that some models have been optimized for the very metrics we employ in AutoRank, either during training or at inference time Freitag et al. (2022a); Finkelstein and Freitag (2024). This can result in artificially inflated scores that do not accurately reflect a model’s true capabilities Kovacs et al. (2024). To mitigate this issue, we aggregate the assessments from multiple learned metrics and LLM-as-a-judge approaches. However, even this strategy has shortcomings. First, scores from different learned metrics often exhibit high correlation among themselves. Second, LLM-as-a-judge approaches, including the Gemba-ESA we use, may also have been utilized to optimize machine translation models.
Another limitation is that we use automatic metrics to evaluate entire paragraphs, whereas their reliability is typically established at the sentence level. Additionally, learned metrics struggle when evaluating translation directions involving low-resource languages, such as English-to-Bhojpuri and English-to-Maasai. Therefore, we evaluate these language pairs using chrF++. However, chrF++ is a surface-level metric that, like BLEU, has been repeatedly shown to correlate poorly with human judgments (Kocmi et al., 2021; Freitag et al., 2022b, 2023).
@@
Given these issues, along with the well-documented biases and limitations of automatic metrics (Karpinska et al., 2022; Moghe et al., 2024), human evaluation remains indispensable. Therefore, the results from human assessments will supersede the automatic rankings presented here.
@@ Appendix A Metrics correlations
The results in appendix˜A show that GEMBA-ESA on CmdA and GPT-4.1 exhibit the highest correlations for almost all languages. In contrast, the weakest correlations are generally observed between xComet and both GEMBA-ESA variants.

### v2

@@ Introduction
We present the preliminary rankings of machine translation (MT) systems submitted to the WMT25 General Machine Translation Shared Task,111www2.statmt.org/wmt25/translation-task.html as determined by automatic evaluation metrics. Because these rankings are derived from automatic evaluation, they may exhibit a bias toward systems that employ re-ranking techniques, such as Quality Estimation or Minimum Bayes Risk decoding. The official WMT25 ranking will be based on human evaluation, which is more reliable and will supersede these results. The official WMT25 ranking will be based on human evaluation, which is more reliable and will supersede these results. The purpose of releasing these findings now is to assist task participants with their system description papers; not to provide final findings.
@@
  • Constrained systems: must use only publicly available training data and models, be limited to a maximum of 20B parameters, and have their model weights released under an open license.
  • Unconstrained systems: (marked with gray) are all other systems. They have no restrictions on training data or model size, and there is no requirement to publish the model weights. This category also includes systems for which training information is not public.
Model # Params Open?
Constrained systems
[…]
Unconstrained systems
[…]
@@
Models.
Our evaluation includes systems submitted by participants, as well as open-weight and proprietary models. We selected the largest or best-performing version of each model where applicable. All constrained and unconstrained systems are listed in Table 1. Full details for all systems will be available in the upcoming WMT25 finding paper.
Prompts.
We prompt all language models using a zero-shot, instruction-following approach, with the specific instructions provided as part of the blind test set. Each model was first tasked with translating the entire document. If this initial attempt failed (e.g., due to producing an incorrect paragraph count or exceeding the token limit), we implemented a fallback strategy of translating the document paragraph by paragraph. This generic setup may disadvantage systems tuned for specific MT instructions, such as TowerLLM or EuroLLM; these are marked with [M].
[…]
Evaluation Data
Languages.
The evaluation covers 32 language pairs, with each test set containing approximately 37K words [^2] As per whitespace split. organized into documents. Half of these language pairs are designated for human evaluation, while the other half belong to a multilingual subtrack evaluated solely with automatic metrics. Most pairs are in the English-to-X direction with the exception of Czech-Ukrainian, Czech-German, and Japanese-Chinese.
Domains.
The test sets combine material from four distinct domains, though language pairs with a non-English source may differ slightly in their domain distribution and dataset size. The domains include:
@@
Data preprocessing.
Each document is divided into segments of approximately 100 words. These segments typically correspond to natural paragraphs; however, some paragraphs were split (or merged) to adhere to this length constraint. Sentence splitting was not performed on the source texts, so segments often contain multiple sentences. For the multimodal context, participants were allowed to (optionally) use accompanying image and video modalities when available, and human annotators who provided reference translations also had access to this context.
All data for this task, including references, system outputs, automatic scores, and the LaTeX source of this document, are publicly available at our repository: github.com/wmt-conference/wmt25-general-mt.
@@
This section details our automatic ranking method, which we refer to as AutoRank. Both the set of automatic metrics and the aggregation procedure have been slightly updated since last year’s shared task.
Metrics.
For most language pairs, [^3] See “Low-resource exception” below. the AutoRank is a combination of three distinct families of evaluation methods:
  • LLM-as-a-Judge (reference-less). We use GEMBA-ESA (Kocmi and Federmann, 2023) with two independent judges: GPT-4.1 (OpenAI, 2025) and Command A (Cohere Team, 2025), both in a reference-less setting.
  • Trained reference-based metrics. Two supervised metrics trained to approximate human quality judgments with references: MetricX-24-Hybrid-XL (Juraska et al., 2024) and XCOMET-XL (Guerreiro et al., 2024).
  • Trained Quality Estimation (QE). The reference-less QE metric CometKiwi-XL (Rei et al., 2023), which is also trained to mimic human judgments.
This combination of reference-based and reference-less (or QE) methods is designed to balance their complementary failure modes. Reference-based metrics typically achieve a higher correlation with human judgments when high-quality references are available, while reference-less methods reduce susceptibility to reference bias when references are suboptimal (Freitag et al., 2023). We also account for known issues with specific metrics. To mitigate a common QE pitfall, i.e., being fooled by fluent output in the wrong language, the GEMBA-ESA prompt explicitly specifies the target language. However, while GEMBA-ESA is intended to reduce bias toward systems that use re-ranking, we note that some participants incorporated it directly as a reward model.
System-level scores.
The system-level score for each language pair is the average of its paragraph-level (segment-level) scores from each metric across the testset. We make one exception for language pairs without human references by excluding CometKiwi-XL from the AutoRank computation. This avoids redundancy, as the other hybrid metrics (MetricX-24-Hybrid-XL and XCOMET-XL) can also run in a reference-less (QE) mode to provide the necessary QE signal.
@@
For the two lowest-resource languages in the testset, i.e., Bhojpuri and Maasai, we rely solely on chrF++ (Popović, 2017), computed with sacrebleu (Post, 2018). This approach was chosen because the reliability of our main metrics is unestablished for these languages (Falcão et al., 2024; Singh et al., 2024; Wang et al., 2024; Sindhujan et al., 2025), whereas high-quality human references required for chrF++ were available.
From system-level scores to AutoRank.
@@
This year’s shared task saw a record number of participants, with submissions from 36 unique teams. [^4] A total of 43 teams initially registered, but 7 later withdrew or were disqualified. Due to budget constraints, we could not include all submissions in the human evaluation. Therefore, we selected a subset of systems for manual evaluation using the Error Span Annotation (ESA) protocol (Kocmi et al., 2024). This subset typically consists of 18 systems per language pair, although this number is higher for some. For any system not selected for human evaluation, the automatic metric ranking (AutoRank) serves as the official final result.
The selection process for human evaluation prioritized constrained systems over unconstrained ones, following a two-step procedure:
@@
    First, the top-8 performing constrained systems were selected, based on their automatic scores.
@@
    Second, the overall top-performing systems from the remaining pool (both constrained and unconstrained) were added until a total of 18 systems was reached.
@@
A key limitation of our evaluation is that some models have been optimized for the very metrics we employ in AutoRank, either during training or at inference time Freitag et al. (2022a); Finkelstein and Freitag (2024). This can lead to artificially inflated scores that do not accurately reflect a model’s true capabilities Kovacs et al. (2024). To mitigate this issue, we aggregate the assessments from multiple learned metrics and LLM-as-a-judge approaches. However, even this strategy has shortcomings. First, scores from different learned metrics often exhibit high correlation among themselves. Second, LLM-as-a-judge approaches, including Gemba-ESA, may themselves have been leveraged to optimize machine translation models.
Another limitation is that we use automatic metrics to evaluate entire paragraphs, whereas their reliability is typically established at the sentence level. Additionally, learned metrics struggle with low-resource languages, such as English-to-Bhojpuri and English-to-Maasai. For these cases, we rely on chrF++ instead. However, chrF++ is a surface-level metric that, like BLEU, has repeatedly been shown to correlate poorly with human judgments (Kocmi et al., 2021; Freitag et al., 2022b, 2023).
@@
Given these challenges, along with the well-documented biases and limitations of automatic metrics (Karpinska et al., 2022; Moghe et al., 2024), human evaluation remains indispensable. Accordingly, results from human assessments will supersede the automatic rankings presented here.
@@ Appendix A Metrics correlations
The results show that GEMBA-ESA on CmdA and GPT-4.1 exhibit the highest correlations for almost all languages. In contrast, the weakest correlations are generally observed between xComet and both GEMBA-ESA variants.